# Complementarity in Human-AI Collaboration: Concept, Sources, and Evidence

### v1

@@ Abstract
Artificial intelligence (AI) can improve human decision-making in various application areas. Ideally, collaboration between humans and AI should lead to complementary team performance (CTP)—a level of performance that neither of them can attain individually. So far, however, CTP has rarely been observed, suggesting an insufficient understanding of the complementary constituents in human-AI collaboration that can contribute to CTP in decision-making. This work establishes a holistic theoretical foundation for understanding and developing human-AI complementarity. We conceptualize complementarity by introducing and formalizing the notion of complementarity potential and its realization. Moreover, we identify and outline sources that explain CTP. We illustrate our conceptualization by applying it in two empirical studies exploring two different sources of complementarity potential. In the first study, we focus on information asymmetry as a source and, in a real estate appraisal use case, demonstrate that humans can leverage unique contextual information to achieve CTP. In the second study, we focus on capability asymmetry as an alternative source, demonstrating how heterogeneous capabilities can help achieve CTP. Our work provides researchers with a theoretical foundation of complementarity in human-AI decision-making and demonstrates that leveraging sources of complementarity potential constitutes a viable pathway toward effective human-AI collaboration.
Keywords Human-AI Complementarity  $\cdot$ Human-AI Decision-Making  $\cdot$ Complementary Team Performance  $\cdot$ Human-AI Teams  $\cdot$ Human-AI Collaboration  $\cdot$ Complementarity Potential
@@ 1 Introduction
The increasing capabilities of artificial intelligence (AI) have paved the way for supporting human decision-making in a wide range of application domains. Examples include decision support for humans in application areas such as customer services (Vassilakopoulou et al.,, 2023), medicine (Jussupow et al.,, 2021), law (Mallari et al.,, 2020), finance (Day et al.,, 2018), and industry (Stauder and Kühl,, 2021). With AI decisions becoming increasingly accurate, there is an obvious temptation to rely on them and to automate decision tasks. However, this approach often falls short of realizing even better performance by combining the unique strengths of the individual members in a human-AI team (Seeber et al.,, 2020). The recent emergence of large language models illustrates this (Malone et al.,, 2023): While applications like ChatGPT often provide helpful, but not always correct results, a human decision-maker can collaborate with the system to, for example, override erroneous answers, in order to achieve superior task performance (Malone et al.,, 2023). Similarly, in the medical domain, both AI models and physicians are able to produce diagnoses individually. It has, however, been demonstrated that humans and AI models could make different errors (Geirhos et al.,, 2021; Steyvers et al.,, 2022). In this context, the AI model might detect patterns in large amounts of data that humans might not discover easily, while humans might excel at the causal interpretation and intuition required to interpret these patterns (Lake et al.,, 2017; Li et al.,, 2019).
This potential for complementarity has inspired researchers to investigate how humans and AI’s individual capabilities could be leveraged to achieve superior team performance instead of either performing the decision task independently. Such an outcome is defined as complementary team performance ($CTP$) (Bansal et al.,, 2021).
Various studies have demonstrated that human-AI teams could outperform human individuals, but often do not exceed the AI model’s individual performance (Bansal et al.,, 2021; Hemmer et al.,, 2021). The observation that $CTP$ is often not attained in these studies raises questions about the factors that contribute to such an outcome. Further, it illustrates that the current knowledge of how to leverage the capabilities of humans and AI for joint decision-making synergies has not yet been sufficiently developed, and that there is a need for additional concepts to foster an in-depth understanding of complementarity in human-AI decision-making. In this work, we therefore pursue the following two research questions:
@@
RQ1: How can we model human-AI decision-making to enable a more nuanced understanding of the synergetic potential in a human-AI team?
@@
We address these research questions by developing a conceptualization of human-AI complementarity that introduces the notion of complementarity potential ($CP$). The conceptualization comprises a formalization of this potential and a delineation of relevant sources. In detail, we argue that complementarity potential has an inherent and a collaborative component. Whereas the first encompasses effects based on different information or capabilities within the human-AI team, the second captures decision-making synergies that only emerge through human-AI interaction. We demonstrate the proposed conceptualization’s application and value in two experimental studies, in which we investigate the effect of two sources of complementarity potential on team performance—the asymmetric distributions of information and capabilities within the human-AI team. In both studies, humans collaborate with an AI model to conduct decision-making tasks. The AI model provides independent decision suggestions that humans can incorporate into their judgment to derive a final team decision. In the first study, we choose the domain of real estate valuation to investigate information asymmetry. We train an AI model to predict real estate prices based on tabular data. Humans receive its suggestions and also have access to a photograph of the real estate. They can use both to make a final team decision. In the second study, we select the context of image classification to analyze the impact of asymmetric capabilities between humans and AI. We train two AI models whose capability gaps differ from those of the human decision-maker. In both studies, we apply our conceptualization to demonstrate that both of these sources increase the inherent complementarity potential and allow $CTP$ to be attained.
@@
To summarize, we make the following contributions to the state of knowledge in information system research: First, we conceptualize human-AI complementarity as a means to comprehensively analyze and design human-AI collaboration. Second, we scrutinize information and capability asymmetries as sources of complementarity potential. Third, we demonstrate the value and application of our concepts and the sources’ potential impact in the two behavioral experiments. This should provide IS researchers with a better understanding and with methodological support when purposefully designing human-AI collaboration in decision-making for more effective outcomes—thereby supporting the development of hybrid intelligence (Dellermann et al.,, 2019), and advocating the “AI with human” (opposed to “AI vs. human”) perspective in societal debates on the future of work (Huysman,, 2020).
@@
In the remainder of this work, we first outline the relevant background and related work in Section 2. In Section 3, we derive the conceptualization of human-AI complementarity. In Section 4, we illustrate our conceptualization’s utility in two experimental studies. We discuss our results in Section 5, before concluding the work in Section 6. We provide the Appendix of this paper at https://github.com/ptrckhmmr/human-ai-complementarity.
@@ 2 Theoretical Foundations and Related Work
In this section, we elaborate on the key concepts and existing work on the collaboration between humans and AI, human-AI complementarity, as well as information and capability asymmetry.
2.1 Collaboration Between Humans and AI
Many terms are used to describe the collaboration between humans and AI. Common ones are human-AI team (Seeber et al.,, 2020), human-AI collaboration (Vössing et al.,, 2022), and human-AI decision-making (Lai et al.,, 2023). These are interrelated concepts that emphasize combining the capabilities of humans and AI. The notion of a human-AI team refers to an organizational setup in which AI agents [^1] Throughout this paper, we refer to agents or systems using machine learning as “Artificial Intelligence” (AI). While we acknowledge the technical distinction between AI and Machine Learning (ML) as discussed, e.g., by Kühl et al., (2022), and although we have considered the more precise reference to an “ML” agent, we adopt the use of the broader “AI” term as the prevalent terminology established in the Computer Science and Human-Computer Interaction communities. This choice reflects the contemporary linguistic trend rather than a lack of distinction between the two fields. or systems are increasingly considered team members rather than support tools for humans—since AI can perform a continuously growing number of tasks independently (Endsley,, 2023; Seeber et al.,, 2020). Human-AI collaboration is the process in which these teams work together in a synergistic manner to achieve shared goals, e.g., with the AI providing recommendations or insights, and humans guiding and refining the AI-generated outputs (Terveen,, 1995; Vössing et al.,, 2022). Human-AI decision-making specifically refers to the collaboration of humans and AI in decision-making tasks. For example, the AI could offer data-driven decision recommendations, while humans leverage their domain expertise, emotional intelligence, and ethical considerations to combine AI recommendations with their judgment to reach a final team decision (Lai et al.,, 2023).
In this article, we focus on human-AI decision-making as a key application area of human-AI collaboration. The increasing capabilities of AI have contributed to its use in a growing number of applications domains, such as medicine (McKinney et al.,, 2020), finance (Kleinberg et al.,, 2018), and customer services (Vassilakopoulou et al.,, 2023). Consequently, AI-based technologies are employed in processes and systems with varying degrees of human involvement, ranging from autonomous decision-making (Rinta-Kahila et al.,, 2022) to auxiliary support for the humans who make the final decision (Bansal et al.,, 2021; Buçinca et al.,, 2020; Lai et al.,, 2020; Liu et al.,, 2021). In this context, unintended or unfair outcomes, e.g., AI-based systems’ decisions that benefit certain individuals more than others (Kordzadeh and Ghasemaghaei,, 2022), have ignited a debate on the degree of autonomy granted to AI to ensure responsible outcomes (Mikalef et al.,, 2022). To alleviate possible detrimental effects, configurations have been suggested that keep humans in the decision-making loop (Grønsund and Aanestad,, 2020; Mikalef et al.,, 2022). Consequently, an increasing number of studies have conducted behavioral experiments to understand how humans make decisions in human-AI teams (Alufaisan et al.,, 2021; Bansal et al.,, 2021; Buçinca et al.,, 2020; Carton et al.,, 2020; Fügener et al., 2021b, ; Fügener et al.,, 2022; Lai et al.,, 2020; Liu et al.,, 2021; Malone et al.,, 2023; Reverberi et al.,, 2022; van der Waa et al.,, 2021; Zhang et al.,, 2022, 2020). In this context, a growing body of work examines how human reliance on AI decisions can be appropriately calibrated to ensure effective decision-making (Buçinca et al.,, 2020; He et al.,, 2023; Kunkel et al.,, 2019; Yu et al.,, 2019; Zhang et al.,, 2020). In application domains comprising high-stakes decisions, e.g., medicine, it is crucial for human experts to identify the AI model’s incorrect suggestions (Jussupow et al.,, 2021). Humans might be helped to judge the AI model’s decision quality by receiving information about the decision’s uncertainty (Fügener et al., 2021a, ; Zhang et al.,, 2020) or by receiving explanations that shed light on the AI model’s decision-making rationale (Bauer et al., 2023a, ).
In order to do so, explainable artificial intelligence (XAI) research has developed various approaches that aim to enable humans to understand the underlying mechanisms that contribute to an AI model’s decision (Adadi and Berrada,, 2018; Bauer et al.,, 2021). Several studies investigate how human-AI decision-making benefits from different XAI techniques. Examples of these range from feature-based techniques (Ribeiro et al.,, 2016) to example-based (van der Waa et al.,, 2021) and rule-based ones (Ribeiro et al.,, 2018). Although experiments have demonstrated the benefits of explanations (Buçinca et al.,, 2020), there is also evidence that explanations might convince humans to follow incorrect AI decisions (Bansal et al.,, 2021) or could foster humans’ propensity to delegate authority to AI by blindly approving its suggestions (Bauer et al., 2023b, ).
A closer look at quantitative studies on human-AI decision-making reveals that, in general, human performance increases when supported by high-performing AI models. In the vast majority of cases, however, the team performance remains inferior to that of the AI model performing the task alone (Hemmer et al.,, 2021; Malone et al.,, 2023). This observation is also in line with the combination of human and technical decisions that the forecasting literature has historically discussed as judgmental adjustments. In this context, several studies find that humans struggle to adjust statistical forecasts effectively (Khosrowabadi et al.,, 2022): Goodwin and Fildes, (1999) show that decision-makers often adjust highly reliable forecasts, while leaving those needing manual intervention untouched. Moreover, Lim and O’Connor, (1995) find that forecasters tend to underestimate statistical forecasts in favor of their judgments. In summary, there are many cases where human adjustments to forecasts and AI suggestions are not beneficial. However, a noteworthy scenario in which the combination of human expertise with technical forecasts could improve the forecast accuracy is in the presence of human contextual knowledge that can serve as a beneficial input for a combined forecast (Blattberg and Hoch,, 1990; Lawrence et al.,, 2006; Sanders and Ritzman,, 1995).
2.2 Human-AI Complementarity
Complementarity between humans and AI is discussed as part of several closely related paradigms: intelligence augmentation, human-machine symbiosis, and hybrid intelligence. Intelligence augmentation is defined as “enhancing and elevating human’s ability, intelligence, and performance with the help of information technology” (Zhou et al.,, 2021, p. 245). It is a form of human-AI collaboration pursuing the idea that machines use their capabilities to assist humans, not necessarily to achieve $CTP$, but to improve human objectives. Human-machine symbiosis is a paradigm that envisions deepening the collaborative connection between humans and AI. It is based on the notion of a symbiotic relationship between both and considers them as a common system rather than two separate entities with the aim to become more effective together compared to working separately (Licklider,, 1960). It also makes the assumption that both entities can offer different capabilities that can be leveraged to overcome human restrictions and to reduce the time needed to solve problems (Gerber et al.,, 2020; Jain et al.,, 2021). Hybrid intelligence pursues the idea of combining human and artificial team members in the form of a socio-technical ensemble. We refer to the work of Dellermann et al., (2019, p. 640), who define hybrid intelligence as “the ability to achieve complex goals by combining human and artificial intelligence, thereby reaching superior results to those each of them could have accomplished separately, and continuously improve by learning from each other.” In addition to realizing performance synergies on an individual level, IS research has also considered AI as a performance driver on an organizational level, when firms develop an AI capability within the organization (Mikalef and Gupta,, 2021). Nevertheless, existing studies under these labels do not provide theoretical views of human-AI complementarity. Articles by Donahue et al., (2022), Steyvers et al., (2022), and Rastogi et al., (2023) are the only works to theorize about human-AI complementarity. Donahue et al., (2022) discuss scenarios in which $CTP$ could occur by considering fairness aspects. Steyvers et al., (2022) derive a framework for combining individual decisions and different types of confidence scores from humans and AI models. Lastly, Rastogi et al., (2023) propose a taxonomy of human and AI strengths together with the notion of across- and within-instance complementarity. All three differ from our approach as they neither investigate different components of complementarity potential or effects, nor do they empirically analyze human-AI decision-making in behavioral experiments.
2.3 Information and Capability Asymmetry
In many application domains, possible sources of complementarity potential could reside in the information and capability asymmetry between humans and AI (Dellermann et al.,, 2019; Ibrahim et al.,, 2021). In this context, it is noteworthy that the mere existence of such asymmetries does not necessarily lead to performance synergies, as it could also have detrimental effects (Dougherty,, 1992).
@@ 2.3 Information and Capability Asymmetry
While AI requires digitally available and sufficient training data to detect patterns that could subsequently be used for decision recommendations (LeCun et al.,, 2015), relevant information might not have been digitized for technical or economic reasons (Ibrahim et al.,, 2021). Contextual information might also not be available to the degree required for it to be included in the AI model training data. Humans, however, could use their expertise by also considering non-digitized information and information about rare events to create a holistic picture in decision-making situations. This might lead to performance synergies when collaborating with AI. This hypothesis is also supported by forecasting theory (Sanders and Ritzman,, 1991, 1995, 2001). Sanders and Ritzman, (1995) analyze the effects of combining statistical forecasts’ predictions with that of a human with access to contextual information, finding that such combinations could impact forecast accuracy positively. Conversely, AI models could also utilize unique contextual information when analyzing information not available to humans in particular decision-making situations. For example, in driving assistance, AI models might infer trajectory suggestions from sensor data to which the driver does not have access.
@@
Research on human teams extensively investigated team composition’s impact on performance (Horwitz,, 2005): On the one hand, asymmetric capabilities of different team members could foster performance synergies, e.g., when these lead to more comprehensive considerations that affect decision-making processes positively (Simons et al.,, 1999). On the other hand, capability asymmetry could also affect team performance negatively due to, e.g., difficulties with developing a shared understanding (Dougherty,, 1992). Similarly, there are also asymmetric capabilities in human-AI teams. Nevertheless, to date the effect of different human and AI capabilities on team performance has remained largely unexplored. Asymmetric capabilities could, e.g., emerge from different internal processing modes. Whereas AI models mostly encode statistical relationships, humans draw upon compositional mental models that encode sophisticated beliefs about the physical and social world (Lake et al.,, 2017; Rastogi et al.,, 2023). While AI models require vast amounts of data for training, humans are able to make inferences from very few data instances (Gopnik and Wellman,, 2012; Lake et al.,, 2017; Tenenbaum et al.,, 2011). In contrast, AI’s internal processing exhibits greater capability to perceive small variations in data than humans can (Findling and Wyart,, 2021). In addition to different processing modes, the nature of the “experiences” that humans and AI are exposed to could also result in asymmetric capabilities. Even though AI models are trained on large datasets, they are restricted to a limited amount of data as input. Humans, on the contrary, accumulate experiences over their entire lifetime, including information across many domains (Dellermann et al.,, 2019; Rastogi et al.,, 2023).

### v2

@@ Abstract
Artificial intelligence (AI) has the potential to significantly enhance human performance across various domains. Ideally, collaboration between humans and AI should result in complementary team performance (CTP)—a level of performance that neither of them can attain individually. So far, however, CTP has rarely been observed, suggesting an insufficient understanding of the principle and the application of complementarity. Therefore, we develop a general concept of complementarity and formalize its theoretical potential as well as the actual realized effect in decision-making situations. Moreover, we identify information and capability asymmetry as the two key sources of complementarity. Finally, we illustrate the impact of each source on complementarity potential and effect in two empirical studies. Our work provides researchers with a comprehensive theoretical foundation of human-AI complementarity in decision-making and demonstrates that leveraging these sources constitutes a viable pathway towards designing effective human-AI collaboration, i.e., the realization of CTP.
Keywords Human-AI Complementarity  $\cdot$ Human-AI Decision-Making  $\cdot$ Complementary Team Performance  $\cdot$ Human-AI Teams  $\cdot$ Human-AI Collaboration  $\cdot$ Complementarity Potential  $\cdot$ Complementarity Effect  $\cdot$ Information Asymmetry  $\cdot$ Capability Asymmetry
@@ 1 Introduction
The increasing capabilities of artificial intelligence (AI) have paved the way for collaborating with humans and supporting them in a wide range of domains. Examples include decision support for humans in application areas such as customer services (Vassilakopoulou et al., 2023), medicine (Jussupow et al., 2021), law (Mallari et al., 2020), finance (Day et al., 2018), and industry (Stauder and Kühl, 2022). With AI decisions becoming increasingly accurate, there is an obvious temptation to fully rely on them and to automate decision tasks. However, this approach often falls short of realizing even better performance by combining and integrating the unique strengths of the individual members in a human-AI team (Seeber et al., 2020). The recent emergence of large language models illustrates this (Malone et al., 2023): While applications like ChatGPT often provide helpful, but not always correct results, a human decision-maker can collaborate with the system to, for example, override erroneous responses in order to achieve superior task performance (Malone et al., 2023). Similarly, in the medical domain, both AI models and physicians are able to produce diagnoses individually. It has, however, been demonstrated that humans and AI models could make different errors (Geirhos et al., 2021; Steyvers et al., 2022) so that they may realize superior results when “teamed up”: For instance, the AI model might detect patterns in large amounts of data that humans might not discover easily, while humans might excel at the causal interpretation and intuition required to understand these patterns (Lake et al., 2017).
This complementarity as the “quality of being different but useful when combined” (Dictionary, 2024) has inspired researchers to investigate how humans’ and AI’s individual abilities could be leveraged to achieve superior team performance compared to either one performing the decision task independently. Such an outcome is defined as complementary team performance (CTP) (Bansal et al., 2021). This phenomenon is of increasing interest as recent studies tell only part of the story: On the one hand, various studies have demonstrated that human-AI teams are able to outperform human individuals (Alufaisan et al., 2021; Inkpen et al., 2023; Liu et al., 2021; Sarkar et al., 2023)—often not analyzing, though, whether they also surpass the AI model’s individual performance (Bansal et al., 2021). On the other hand, while many settings exist in which humans still show better task performance (Grace et al., 2018, 2024), recent studies also demonstrate evidence for human-AI teams outperforming AI models (Dvijotham et al., 2023; Fügener et al., 2022; Ma et al., 2023).
To make things even more complex, the particular design of human-AI collaboration in decision-making affects humans and their contributions within the team, e.g., the use of unique human knowledge (Fügener et al., 2021), the self-assessment of human capabilities (Fügener et al., 2022), the adjustments of mental models (Bauer et al., 2023), the incentive for “active rethinking” (Lu and Zhang, 2024), or human task performance and learning (Förster et al., 2024). In summary, current research still misses compelling explanations for the success of human-AI teams as well as a systematic understanding of complementarity when the team performance is measured against the performance of both team members individually. Thus, concentrating on decision-making as an important application of human-AI collaboration (Lai et al., 2023), we pursue the following two research questions in this work:
@@
RQ1: How can we model human-AI collaboration in decision-making to enable a more nuanced understanding of the synergetic potential in a human-AI team?
@@
We address these research questions by developing a conceptualization of human-AI complementarity that introduces and quantifies complementarity potential ($CP$) and complementarity effect ($CE$) and outlines the two key sources of complementarity—information and capability asymmetry. In detail, we argue that both complementarity potential and complementarity effect consist of an inherent and a collaborative component. Whereas the first component captures decision-making synergies that, for each task instance, can be attributed to the individually more accurate team member within the human-AI team, the second component captures decision-making synergies that only emerge through collaboration resulting in team decisions which are more accurate than each of the individual ones. We demonstrate the application and the value of our conceptualization in two experimental studies—leveraging the two sources of complementarity, i.e., information and capability asymmetry within the human-AI team. In both studies, humans collaborate with an AI model to conduct decision-making tasks. The AI model provides independent decision suggestions that humans can incorporate into their judgment to derive a final team decision. In the first study, we choose the domain of real estate valuation to investigate information asymmetry. We train an AI model to predict real estate prices based on tabular data. Humans receive its suggestions and also have access to a photograph of the real estate. They can use both to arrive at a final team decision. In the second study, we select the context of image classification to analyze the impact of capability asymmetry between humans and AI. We train two AI models whose capability gaps differ from those of the human decision-maker. In both studies, we apply our conceptualization to demonstrate that both of these sources increase the inherent component of complementarity potential as well as the realized effect, resulting in CTP.
@@
To summarize, we make the following contributions to the state of knowledge in information systems (IS) research: First, we conceptualize human-AI complementarity as a means to comprehensively analyze and design human-AI collaboration in decision-making. Second, we scrutinize information and capability asymmetries as sources of complementarity. Third, we demonstrate the value and application of our concepts and the sources’ potential impact in two behavioral experiments. This should provide IS researchers with a better understanding and methodological support when purposefully designing human-AI collaboration in decision-making for more effective outcomes—thereby supporting the development of hybrid intelligence (Dellermann et al., 2019) and advocating the “AI with human” (opposed to “AI vs. human”) perspective in societal debates on the future of work (Huysman, 2020).
@@
In the remainder of this work, we first outline the relevant background and related work in Section 2. In Section 3, we derive the conceptualization of human-AI complementarity. In Section 4, we empirically illustrate our conceptualization’s utility in two experimental studies. We discuss our results in Section 5, before concluding the work in Section 6. We provide the Appendix of this paper at https://github.com/ptrckhmmr/human-ai-complementarity.
@@ 2 Theoretical Foundations and Related Work
In this section, we elaborate on the key concepts and on existing work that provide the foundation for a formal conceptualization of complementarity. We first cover existing differences between humans and AI models [^1] Throughout this paper, we refer to models or systems using machine learning as “Artificial Intelligence” (AI) (Berente et al., 2021; Collins et al., 2021; Rai et al., 2019). While we acknowledge the technical distinction between AI and Machine Learning (ML) as discussed, e.g., by Kühl, Schemmer, Goutier and Satzger (2022) and although we have considered the more precise reference to an “ML” model, we adopt the use of the broader “AI” term as the prevalent terminology established in the Computer Science and Human-Computer Interaction communities. This choice reflects the contemporary linguistic trend rather than a lack of distinction between the two fields. as sources of complementarity, then review existing literature on human-AI collaboration in decision-making as the means to harness complementarity, and finally summarize the current state of knowledge on human-AI complementarity.
2.1 Sources of Complementarity: Information and Capability Asymmetry
In the following, we assume that both human and AI can independently produce solutions to a particular decision problem. Collaboration would not be of interest if both always generated identical decisions. However, human and AI typically make different types of errors (Geirhos et al., 2021; Steyvers et al., 2022)—a phenomenon that very generally can be traced back to two key sources. To illustrate this, we employ a simple Input-Processing-Output Model for decision-making (Figure 1)—as used both in IS (D’Arcy et al., 2014; Parsons and Wand, 2012) and social psychology (Gladstein, 1984; Hackman and Morris, 1975): Any team player (human or AI) draws on and processes certain sets of information to derive a decision. Discrepancies in the decision outcome may either stem from different levels of available information (information asymmetry) or from different capabilities to process this information (capability asymmetry). The two asymmetries may then be exploited for a (hopefully) improved team decision resulting in CTP.
Figure 1: Asymmetries in decision-making between human and AI as sources of complementarity.
@@ 2.1 Sources of Complementarity: Information and Capability Asymmetry
Often, humans and AI have access to different sets of information as decision input: On the one hand, AI models are trained on a well-defined, limited, and digitally available set of data (LeCun et al., 2015). Humans, however, may also have access to information that—due to technical or economic reasons—is not digitized and, thus, not usable for an AI model (Ibrahim et al., 2021). They may use contextual information or information on rare events for a more holistic decision-making setting. For example, AI models in the medical domain derive diagnoses from X-ray scans by analyzing the pixels of the image (Irvin et al., 2019). Human radiologists, however, may also draw on information from direct interaction with patients or from access to medical records. On the other hand, the AI may also have access to information not available to the human decision-maker: A driving assistance system may base its actions on sensor data, e.g., from lidar systems that are not accessible by the human while driving (Li and Ibanez-Guzman, 2020).
@@
Even if the information would be identical, different outcomes may be driven by different internal modes of information processing between human and AI (Kühl, Goutier, Baier, Wolff and Martin, 2022). Such capability asymmetries emerge as AI models encode relationships inferred from training data, whereas humans employ compositional mental models that encode beliefs about the physical and social world (Lake et al., 2017; Rastogi et al., 2023). While humans are capable of conducting decision-making tasks already after a few trials, AI models require vast amounts of training data to become capable of accurately performing decision-making tasks (Gopnik and Wellman, 2012; Kühl, Goutier, Baier, Wolff and Martin, 2022; Lake et al., 2017; Tenenbaum et al., 2011). Similarly, humans have often gained experiences with regard to a particular decision-making task continuously over their lifetime, while AI models’ experiences are limited to the task instances seen during model training and may only be updated from time to time (Dellermann et al., 2019; Rastogi et al., 2023). On the other hand, AI may excel in digesting vast amounts of information much faster than humans could, and its processing exhibits a greater capability to perceive even small variations in data (Findling and Wyart, 2021).
Overall, both asymmetries may give rise to performance synergies when collaborating in human-AI teams. A traditional example is forecasting theory (Sanders and Ritzman, 1991, 1995, 2001): Sanders and Ritzman (1995) analyze the effects of combining statistical forecasts with predictions of a human with access to contextual information and find that such combinations could positively impact forecast accuracy. Composing suitable teams to optimize team performance has been investigated for human-only teams (Horwitz, 2005), but lately has also been applied to human-AI teams (Hemmer et al., 2022). While research has shown that such asymmetries can improve team performance (Simons et al., 1999; Zhang et al., 2022), they also bear the risk of negative effects, i.e., performance degradation (Dougherty, 1992): Heterogeneous and interdisciplinary human-only teams may enjoy benefits from higher levels of individual problem-solving creativity that may be outweighed, though, by the difficulties to effectively communicate with each other (Ancona and Caldwell, 1992).
2.2 Means for Harnessing Complementarity: Human-AI Collaboration
Many terms are used to describe the interplay between humans and AI. Common ones are human-AI team (Seeber et al., 2020), human-AI collaboration (Vössing et al., 2022), and human-AI decision-making (Lai et al., 2023). These are interrelated concepts that emphasize combining the complementary qualities of humans and AI. The notion of a human-AI team refers to an organizational setup in which AI systems are increasingly considered equitable team members rather than support tools for humans—since AI can perform a continuously growing number of tasks independently (Endsley, 2023; Seeber et al., 2020). Human-AI collaboration is the process in which these teams work together in a synergistic manner to achieve shared goals, e.g., with the AI providing recommendations or insights, and humans guiding and refining the AI-generated outputs (Terveen, 1995; Vössing et al., 2022). Human-AI decision-making refers specifically to human-AI collaboration in decision-making tasks. For example, the AI could offer data-driven decision recommendations, while humans leverage their domain expertise, emotional intelligence, and ethical considerations to combine AI recommendations with their judgment to reach a final team decision (Lai et al., 2023).
In this article, we focus on human-AI decision-making as a key application area of human-AI collaboration. The increasing abilities of AI have contributed to its use in a growing number of application domains, such as medicine (McKinney et al., 2020), finance (Kleinberg et al., 2018), and customer services (Vassilakopoulou et al., 2023). Consequently, AI-based technologies are employed in processes and systems with varying degrees of human involvement, ranging from autonomous decision-making (Rinta-Kahila et al., 2022) to just auxiliary support for the ultimate human decision-maker (Bansal et al., 2021; Buçinca et al., 2020; Lai et al., 2020; Liu et al., 2021). In this context, unintended or unfair outcomes, e.g., AI-based systems’ decisions that benefit certain individuals more than others (Kordzadeh and Ghasemaghaei, 2022), have ignited a debate on the degree of autonomy granted to AI to ensure responsible outcomes (Mikalef et al., 2022). To alleviate possible detrimental effects, configurations have been suggested that keep humans in the decision-making loop (Grønsund and Aanestad, 2020; Pathirannehelage et al., 2024; Mikalef et al., 2022).
Researchers have been devoting significant efforts to better understand human-AI decision-making and to design the collaboration in a way that ultimately achieves CTP (Hemmer et al., 2021; Lai et al., 2023). Overall, a wide range of behavioral experiments seek to help us understand how humans make decisions within human-AI teams (Alufaisan et al., 2021; Bansal et al., 2021; Buçinca et al., 2020; Carton et al., 2020; Fügener et al., 2021, 2022; Lai et al., 2020; Liu et al., 2021; Malone et al., 2023; Reverberi et al., 2022; van der Waa et al., 2021; Zhang et al., 2022, 2020). A key emerging concept in this space is that of human reliance on AI advice that needs to be appropriately calibrated to ensure effective decision-making (Buçinca et al., 2020; He et al., 2023; Kunkel et al., 2019; Schemmer et al., 2022; Yu et al., 2019; Zhang et al., 2020). To assist the human in judging the AI’s decision quality, the AI can provide information about the decision’s uncertainty (Fügener et al., 2021; Zhang et al., 2020). Similarly, it may deliver various types of explanations that shed light on its decision-making rationale (Adadi and Berrada, 2018; Bauer et al., 2023), e.g., using feature-based (Ribeiro et al., 2016), example-based (van der Waa et al., 2021) or rule-based techniques (Ribeiro et al., 2018). In application domains involving high-stakes decisions, e.g., medicine, it is crucial for human experts not to exhibit a certain reliance “on average” but in particular to identify the AI model’s incorrect suggestions (Jussupow et al., 2021). Thus, the calibration of reliance needs to reflect the actual decision quality (Schoeffer et al., 2023).
A closer look at quantitative studies on human-AI decision-making reveals that, in general, human performance increases when supported by high-performing AI models. In the vast majority of cases, however, the team performance remains inferior to that of the AI model when performing the task alone (Hemmer et al., 2021; Malone et al., 2023). This means that joint decision-making currently does not lead to the realization of the full complementarity potential: Humans often do not show appropriate reliance by contributing their own decision capabilities in the right places. While recent studies have shown that the performance of human-AI teams can improve beyond that of the individual team members (Dvijotham et al., 2023; Fügener et al., 2022; Ma et al., 2023), the underlying mechanisms why performance synergies often fail to materialize are still poorly understood. Research has explored different paths: First, humans’ ability to exert appropriate reliance depends on the overall decision-making situation, e.g., whether it is possible to ex post verify the correctness of the decisions (de Véricourt and Gurkan, 2023). Second, “imperfections” on the human side may contribute to this: Humans can struggle to correctly assess their own capabilities in comparison to that of the AI (Fügener et al., 2022). They may develop implicit biases against the AI that inhibit their willingness to rely on its advice (Turel and Kalhan, 2023), or they may start to mirror the AI’s behavior by following its advice even when it is incorrect (Fügener et al., 2021). Third, humans may be misled by signals from the AI that were originally intended to help them better assess its decision quality. Additional explanations may not be correct (Morrison et al., 2024), or they may distort humans’ situational balancing of available information, leading to misconceptions and suboptimal decisions (Bauer et al., 2023). Overall, further research is required to provide the means for developing a holistic understanding of the synergetic potential between humans and AI—an objective this work pursues by proposing a conceptualization of human-AI complementarity.
2.3 Concepts Related to Human-AI Complementarity
Complementarity between humans and AI is discussed as part of several closely related paradigms: Intelligence augmentation, human-machine symbiosis, and hybrid intelligence.
Intelligence augmentation is defined as “enhancing and elevating human’s ability, intelligence, and performance with the help of information technology” (Zhou et al., 2021, p. 245). It follows the idea that machines use their abilities to assist humans, not necessarily to achieve CTP, but to improve human objectives. Human-machine symbiosis is a paradigm that envisions deepening the collaborative connection between humans and AI. It is based on the notion of a symbiotic relationship between both and considers them as a common system rather than two separate entities with the aim of becoming more effective together than working separately (Licklider, 1960). It also makes the assumption that both entities can offer different capabilities that can be leveraged to overcome human restrictions and to reduce the time needed to solve problems (Gerber et al., 2020; Jain et al., 2021). Hybrid intelligence pursues the idea of combining human and AI team members in the form of a socio-technical ensemble. We refer to the work of Dellermann et al. (2019, p. 640), who define hybrid intelligence as “the ability to achieve complex goals by combining human and artificial intelligence, thereby reaching superior results to those each of them could have accomplished separately, and continuously improve by learning from each other.” In addition to realizing performance synergies on an individual level, IS research has also considered AI as a performance driver on an organizational level when firms develop an AI capability within the organization (Mikalef and Gupta, 2021).
Nevertheless, existing studies under these labels do not provide theoretical views of human-AI complementarity. Articles by Donahue et al. (2022), Steyvers et al. (2022), and Rastogi et al. (2023) are the only works to theorize about human-AI complementarity. Donahue et al. (2022) discuss scenarios in which CTP could occur by considering fairness aspects. Steyvers et al. (2022) derive a framework for combining individual decisions and different types of confidence scores from humans and AI models. Lastly, Rastogi et al. (2023) propose a taxonomy of human and AI strengths together with the notion of across- and within-instance complementarity. All three differ from our approach as they do not distinguish between complementarity potential ($CP$) and complementarity effect ($CE$) with their respective components. Moreover, they do not empirically analyze human-AI decision-making in behavioral experiments.